# RIDAS: A Multi-Agent Framework for AI-RAN with Representation- and Intention-Driven Agents

### v2

@@ Abstract
Index Terms:
6G, AI-RAN, AI agents, resource allocation, large language models.
@@ III-A2 SVID
As shown in Eq.(5), the reconstruction error of SVID is provably no greater than that of SVD. Furthermore, Proposition 1 can be straightforwardly generalized to $r$ decompositions. Accordingly, SVID not only achieves a lower-error approximation $U_{r}\Sigma_{r}V_{r}^{T}$ than SVD, but also offers a controllable trade-off between representation dimension and distortion, thereby providing a preliminary solution to Challenge 1.
@@ III-B2 Planning
The proposed RDA and IDA together form the RIDAS framework. On the one hand, RDAs expose an interface for controlling their representation actions, thereby satisfying diverse user QoS requirements with minimal resource consumption through the adjustment of the control parameter $\boldsymbol{\theta}$. On the other hand, the IDA interprets user intent to configure the control parameters of RDAs in a manner that further minimizes system resource usage. Consequently, RIDAS introduces a novel AI-RAN paradigm, demonstrating how LLM-based agents can interact with and manage the underlying AI models.
@@ IV-A1 Scenario setup
  • Low: QoS demands accuracy $>$ 70%, corresponding to shorter bit-stream lengths;
  • Medium: QoS requires accuracy $>$ 80%;
  • High: QoS requires accuracy $>$ 90%, corresponding to longer bit-stream lengths.
@@
[…]

### v4

@@ III-A2 SVID
As shown in Eq.(5), the reconstruction error of SVID is provably no greater than that of SVD. Furthermore, Proposition 1 can be straightforwardly generalized to rank-$r$ decompositions. Accordingly, SVID not only achieves a lower-error approximation $U_{r}\Sigma_{r}V_{r}^{T}$ than SVD, but also offers a controllable trade-off between representation dimension and distortion, thereby providing a preliminary solution to Challenge 1. The computational complexity of SVID is dominated by the truncated SVD step, resulting in an overall complexity of $\mathcal{O}(mnr)$.
@@ III-B2 Planning
The proposed RDA and IDA together form the RIDAS framework. On the one hand, RDAs expose an interface for controlling their representation actions, thereby satisfying diverse user QoS requirements with minimal resource consumption by adjusting the control parameter $\boldsymbol{\theta}$. On the other hand, the IDA interprets user intent to configure the control parameters of RDAs in a manner that further minimizes system resource usage. Consequently, RIDAS introduces a novel AI-RAN paradigm, demonstrating how LLM-based agents can interact with and manage the underlying AI models.
@@ IV-A1 Scenario setup
  • Low: QoS demands accuracy >> 70%, corresponding to shorter bit-stream lengths;
  • Medium: QoS requires accuracy >> 80%;
  • High: QoS requires accuracy >> 90%, corresponding to longer bit-stream lengths.
@@
$\text{Bandwidth (MHz)}=\frac{\lvert\mathcal{B}\rvert\,/\,\text{code rate}}{\text{transmission time}}\times\frac{1}{\log_{2}\bigl(1+10^{\frac{\text{SNR}}{10}}\bigr)\times 10^{6}}\,.$ (7)